How do I solve this equation for the value of x?
(3x - 5) / 7 = (3x + 5) / 5

Answer by mananth(16949)

(3x - 5) / 7 = (3x + 5) / 5
5(3x-5)=7(3x+5)
15x-25=21x+35
6x= -60
x= -10